OpenStudy (anonymous):

1/4(2x-14)=4 solve for x

OpenStudy (anonymous):

this is my way of solving it: 1. multiply both sides by the denominator to get rid off the fractions 2. then simplify and solve for "x"

OpenStudy (tkhunny):

Always look for ways to make your life simpler. \(\dfrac{1}{4}(2x-14) = \dfrac{1}{4}(2)(x-7) = \dfrac{1}{2}(x-7) = 4\)